OpenStudy (anonymous):

What is x in this quadratic equation? 3x(squared) -8 + 4x(squared)= 5x(squared) + 48

OpenStudy (anonymous):

\[3x ^{2} - 8 + 4x ^{2} = 5x ^{2} +48\]

Directrix (directrix):

Collect variable terms on the left and constant terms on the right side of the equation.

OpenStudy (anonymous):

I did that, but my friend has a different answer than me

OpenStudy (anonymous):

What is the answer?

Directrix (directrix):

I need to see your work or your friend's work.

Directrix (directrix):

I do need to ask if you are solving for x over the set of Real Numbers or over the set of Complex Numbers.

OpenStudy (anonymous):

real numbers

Directrix (directrix):

Okay, now the work or the final simplified equation you and/or the friend got.

OpenStudy (anonymous):

When i collected variable terms, i ended up with 2x(squared)= 56 then x(squared) = 56/2 x(squared) = 28 x= plus or minus, square root of 28

OpenStudy (anonymous):

I don't know my friend's work, I just know she got a different answer than me

Directrix (directrix):

When I collected like terms and simplified, I got: x^2 = - 28 One of us has a sign error.

Directrix (directrix):

And, I think it is I with the sign error.

OpenStudy (anonymous):

alright, thank you very much

OpenStudy (anonymous):

one more question if you don't mind

Directrix (directrix):

Wait a second. You need to simplify the radicals.

OpenStudy (anonymous):

\[2\sqrt{6}\]

Directrix (directrix):

x= plus or minus, square root of 28 x = ± 2 * √ 7

Directrix (directrix):

28 = 2*2*7

OpenStudy (anonymous):

OOOh okay, thank you!

OpenStudy (anonymous):

I get it now

Directrix (directrix):

What is the other question?

OpenStudy (anonymous):

That was it

Directrix (directrix):

Oh, okay. Better tell the friend. :)

OpenStudy (anonymous):

Haha I did

OpenStudy (anonymous):

Thanks again :D

Directrix (directrix):

Glad to help.